Potential difference

Charge Q=+2.00uC is distributed uniformly over the volume of an insulating sphere that has radius R=6.00cm. What is the potential difference between the center of the sphere and the surface of the spheres?
Answer #1

V at the center = 3/2*k*Q/R


V at the surface = k*Q/R


Potential difference = 0.5*9*10^9*2*10^-6 / 0.06

= 1.5 * 10^5 V


answered by: alida